# A blueprint for large-scale quantum-network deployments

### v1

@@ IV-C Management, administration and operation
In MadQCI, several management, administration and operation (OAM) mechanisms coexisted and were tested. Relevant to this paper is the software-defined management and operation based on ETSI GS QKD standards implemented at UPM, named SD-QKD Stack in this work. It is based on the approach explained in section III-D and depicted in Fig. 3. However, there were also some QKD vendor-specific key and network management systems deployed, as well as the management systems of the classical resources.
@@
The SD-QKD Stack enabled the end-to-end quantum services by consuming the QKD keys generated by the links described in the previous section. It is composed of the SDN controller, the set of its agents and another set of local key management systems (LKMS) to perform the key management and forwarding tasks, according to the ETSI GS QKD 015 specification. This system therefore implements both key management and networking functionalities. Thus, a joint management of both network and cryptographic key resources was achieved, and a QoS-driven service as specified in ETSI GS QKD 004 could be provided [4].
@@
It operates as follows. First, when a service is requested via the ETSI GS QKD 004 interface at two endpoints—or using the ETSI GS QKD 014 ones—, the agents at each node collected the necessary information and relayed it to the controller, that decided which settings are suitable in each case. Then, it commands the agents to configure the resources of each node, both the LKMS and the QKD systems, to provide the service. The needed keys generated by the QKD systems on each link of the calculated path are then collected, independently of their manufacturer, and managed by each LKMS. As this component includes the FM of each node, it consumes the keys to perform the key forwarding along all the required hops in a vendor agnostic way.
@@
In an ideal approach, this operation and management is joint with that of the QKD resources. Specifically, the ETSI GS QKD 015 specification defines some optical-networking related fields in its data model, thereby enabling this integrated software-defined control. For instance, the SDN controller is able to configure the set of SDN agents the required wavelength and expected attenuation in each hop, to configure these requirements in the SDN-aware devices, both optical and QKD. This shows the network programmability that SDN promotes, since those optical-networking requirements may arise from the quality of service requested by a subscriber using the ETSI GS QKD 004 interface, together with all other connectivity and key management requirements.
However, each system has a different degree of maturity, so such integration is an ongoing task and only specific experiments have been performed, as the following section V explains. Another approach is using other interoperability mechanisms, such as the optical-quantum orchestration promoted by the ETSI-GS-QKD-018 specification, that it is also an ongoing task.
@@ V-1 General overview
In general terms, the installation and connection tasks were successful in every domain and had far exceeded the objectives of the associated research projects. Table II gathered the MadQCI main features and capabilities.
@@ V-3 Management, administration and operation
Also, using the same approach, other quantum-enabled services have been trialed, such as quantum oblivious key distribution [58, 8] of quantum digital signature [59]. Of interest for this work is the first case, where some CV-QKD in the co-located domain were configured to deliver the raw key, without post-processing. Other DV-QKD still delivered symmetric distilled key. Using these two resources, key material suitable for quantum oblivious transfer could be obtained. This is an example of the SDN-aware interoperability needed to perform the end-to-end service. The SD-QKD Stack was modified to provide this service using the same software components and, as far as possible, the same standard interfaces. When not, the syntax of the SDN models was overloaded in order to contain the new configurations. Also, the configuration of operating parameters, such as wavelength, was trialed in Huawei CV-QKD systems using the QuAM/QuAI novel technology [60], so that the switching capabilities in the co-located domain could be configured using the same SD-QKD Stack.
@@ VI Conclusion and future work
In this way, it has been possible to present its results as a blueprint for large-scale quantum network deployments. The fundamentals of optical networking and quantum communications to support this goal have also been discussed. It has been important to characterise the networks in production, taking into account their complexity beyond the most basic components.
@@
The quantum connectivity achieved on each link has allowed to show how the SDN operation and management approach can be used to extend the performance to end-to-end quantum-enhanced services. The SD-QKD Stack enabled the testing of a large number of use cases. This software implemented several ETSI-GS-QKD standards, ensuring that it is a valid solution for deployments with a more ambitious scope. It is based on a centralised operation and management of resources through a logically centralized controller and a set of agents at each node, as well a set of local KMS that performs both the key management and networking functionalities. Thus, not only encryption applications have been supported, but also other less tested primitives, such as oblivious transfer.
MadQCI, is committed to be the reference instance in the deployment of other QCI and the pan-European quantum communication infrastructure EuroQCI. In fact, this network is now starting a new iteration thanks to investments related to the Quantum Flagship and Digital Europe Programme, as well as the MADQuantum-CM project [61], part of the Recovery and Resilience Facility funds managed by Spain. Design and procurement is underway and details are being finalised for collaboration with various partners to promote higher value-added use cases, including those involving the satellite segment.
@@ Appendix A Schemes
This appendix gathers the most relevant nodal schemes. These reflect the application of the approach described in the text, as well as other details specific to each node. The nodes not depicted do not contain relevant details and their optics resemble the optics used to illuminate the fibre linking them. The legend used in figure 10 is the same for figures 11, 12, 13 and 14.

### v2

@@ IV-C Operation, administration and management
In MadQCI, several operation, administration and management (OAM) mechanisms coexisted and were tested. Relevant to this paper is the software-defined management and operation based on ETSI GS QKD standards implemented at UPM, named SD-QKD Stack here. It is based on the approach explained in section III-D and depicted in Fig. 3. However, there were also some QKD vendor-specific key and network management systems deployed, as well as the management systems of the classical resources.
@@
The SD-QKD Stack enabled the end-to-end quantum services by consuming the QKD keys generated by the links described in the previous section. It is composed of the SDN controller, the set of its agents at the trusted nodes and a set of local key management systems (LKMS) to perform the key management and forwarding tasks, according to the ETSI GS QKD 015 specification. This system therefore implements both key management and networking functionalities. Thus, a joint management of both network and cryptographic key resources was achieved, and a QoS-driven service as specified in ETSI GS QKD 004 could be provided [4].
@@
It operates as follows. First, when a service is requested via the ETSI GS QKD 004 interface at two endpoints—or using the ETSI GS QKD 014 ones—, the agents at each node collect the necessary information and relay it to the controller, that decided which settings are suitable in each case. Then, it commands the agents to configure the resources of each node, both the LKMS and the QKD systems, to provide the service. The needed keys generated by the QKD systems on each link of the calculated path are then collected, independently of their manufacturer, and managed by each LKMS. As the latter component includse and invokes the FM of each node, it consumes the keys to perform the key forwarding along all the required hops in a vendor agnostic way.
@@
In an ideal approach, this operation and management is combined with that of the QKD resources. Specifically, the ETSI GS QKD 015 specification defines some optical-networking related fields in its data model, thereby enabling integrated software-defined control. For instance, the SDN controller is able to configure the set of SDN agents the required wavelength and expected attenuation in each hop, to configure these requirements in the SDN-aware devices, both optical and QKD. This demonstrates network programmability that SDN promotes, since the optical-networking requirements may arise from the quality of service requested by a subscriber using the ETSI GS QKD 004 interface, together with all other connectivity and key management requirements.
However, each system has a different degree of maturity, so such integration is an ongoing task and only specific experiments have been performed, as the following section V explains. Another approach is using other interoperability mechanisms, such as the optical-quantum orchestration promoted by the ETSI-QKD-GS-018 specification, that it is also an ongoing task.
@@ V-1 General overview
In general terms, the installation and connection tasks were successful in every domain and had far exceeded the objectives of the associated research projects.
@@ V-3 Management, administration and operation
Also, using the same approach, other quantum-enabled services have been trialed, such as quantum oblivious key distribution [61, 9] and quantum digital signature [62]. Of interest for this work is the first case, where some CV-QKD in the co-located domain were configured to deliver the raw key, without post-processing. Other DV-QKD still delivered symmetric distilled key. Using these two resources, key material suitable for quantum oblivious transfer could be obtained. This is an example of the SDN-aware interoperability needed to perform the end-to-end service. The SD-QKD Stack was modified to provide this service using the same software components and, as far as possible, the same standard interfaces. When not, the syntax of the SDN models was overloaded in order to contain the new configurations. Also, the configuration of operating parameters, such as wavelength, was trialed in Huawei CV-QKD systems using the QuAM/QuAI novel technology [63], so that the switching capabilities in the co-located domain could be configured using the same SD-QKD Stack.
@@ VI Conclusion and future work
In this way, it has been possible to present the results as a blueprint for large-scale quantum network deployments. The fundamentals of optical networking and quantum communications to support this goal have also been discussed. It has been important to characterise the networks in production, taking into account their complexity beyond the most basic components.
@@
The quantum connectivity achieved on each link has allowed to show how the SDN operation and management approach can be used to extend the performance to end-to-end quantum-enhanced services. The SD-QKD Stack enabled the testing of a large number of use cases. This software implemented several ETSI-QKD standards, ensuring that it is a valid solution for deployments with a more ambitious scope. It is based on a centralised operation and management of resources through a logically centralized controller and a set of agents at each node, as well as of LKMS modules. Not only encryption applications have been supported, but also other less proliferated primitives, such as, e.g., oblivious transfer.
MadQCI, aims to be the reference instance in the deployment of other QCI and the pan-European quantum communication infrastructure EuroQCI. In fact, this network is now starting a new iteration thanks to investments of the Quantum Flagship and Digital Europe Programme, as well as the MADQuantum-CM project [64], financed by the Recovery and Resilience Facility funds managed by Spain. Design and procurement is underway and details are being finalised for collaboration with various partners to promote higher value-added use cases, including those involving the satellite segment.
@@ Appendix A Schemes
This appendix gathers the most relevant node diagrams, which apply the described approach. The nodes not depicted do not contain additional relevant details and are the mirror image of the equipment that illuminates the fibre linking them. The legend used in figure 11 is the same for figures 12, 13, 14 and 15.